OpenStudy (anonymous):

Find the slope of the tangent line to the polar curve r=2sin(theta) for the given value (theta)=(pi/6).

OpenStudy (anonymous):

\[dy/dx= (rcos \theta+\sin \theta *dr/d \theta)/(-rsin \theta+\cos \theta*dr/d \theta)\] This is the equation we're supposed to use, and when I try it, ut works out to: \[\sqrt{3}+3\] But the answer is supposed to be: \[\sqrt{3}\] I'm not sure what I'm doing wrong becuase I've checked my work and nothing looks wrong.

OpenStudy (tkhunny):

Please show your result for \(\dfrac{dr}{d\theta}\).

OpenStudy (anonymous):

\[dr/d \theta=2\cos\]

OpenStudy (anonymous):

*2cos(theta)

OpenStudy (tkhunny):

Right. I was going to call you out on that! Good work beating me to it.

OpenStudy (anonymous):

Typos, haha. Sorry about that.

OpenStudy (tkhunny):

How did you proceed after that? Did you use 'r' in the giant fraction or \(2\sin(\theta)\)?

OpenStudy (anonymous):

I used 2sin(theta)

OpenStudy (tkhunny):

Okay, so Numerator \(4\sin(\theta)\cos(\theta)\) How did that go?

OpenStudy (anonymous):

How did you get that? I got \[(2\sin \theta)\cos \theta+\sin \theta(2\cos \theta)\]

OpenStudy (anonymous):

nevermind, you simplified, right?

OpenStudy (anonymous):

Then for the denominator \[-(2\sin \theta)\sin \theta+\cos \theta(2\cos \theta)\] ?

OpenStudy (anonymous):

@tkhunny ?

OpenStudy (tkhunny):

Couldn't find the thread. Not sure why. I'm back. That's it. Now, use your best trigonometry to change to functions of \(2\theta\).